# Preference Tuning For Toxicity Mitigation Generalizes Across Languages

### v1

@@ Acknowledgement
Appendix A Training Details
@@ A.1 DPO Preference Tuning
We use HuggingFace trl library and follow Lee et al.’s (2024) hyperparameters (except learning rate) for full model finetuning of mGPT and BLOOM-1.7B. For QLoRA finetuning of Aya-23, LLama2, and Llama3, we apply QLoRA (Dettmers et al., 2023) on each model layer, with a rank of 64, a scaling parameter of 16 and a dropout of 0.05. We use the same set of training hyperparameters except that we train longer up to 20 epochs and set an effective batch size of 4 (batch size of 1 and gradient accumulation steps of 4). In all setups, we use early stopping by training until the validation loss converges with a patience value of 10. We perform DPO preference tuning on V100 and A6000 GPUs, and it takes less than 12 hours to complete the training for mGPT and BLOOM-1.7B and around 24 hours to complete the training for Aya-23, Llama2 and Llama3.
@@
Table 5: Hyperparameters for DPO preference tuning for mGPT and BLOOM (1.7B).
@@ A.2 Probe Training
We train the linear probe $w_{\mathrm{toxic}}$ for English binary toxicity classification with seed 99 on 90% of 159,571 comments from the Jigsaw dataset cjadams et al. (2017). Table 5 displays the hyperparameters used for training. It achieves a validation accuracy of 94.31% on the remaining 10% dataset. In addition, the ROC-AUC (Receiver Operating Characteristic - Area Under the Curve) score on the test split of Jigsaw dataset is 0.862. The whole experiment was conducted on a single NVIDIA RTX A6000 for approximately 50 hours.
@@
Table 6: Training hyperparameters for the binary toxicity classification probe $w_{\mathrm{toxic}}$.
@@
Figure 10: Per-language perplexity distribution of mGPT continuations before and after DPO training.
Appendix B Distribution of Perplexity Scores
@@
Figure 11: Tradeoffs between DPO learning rate, toxicity in post-DPO generation and perplexity across 17 languages.
Appendix C Multilingual Capability Evaluation after DPO
We show little degradation on model’s multilingual capability after DPO preference tuning. In fact, some languages even experience slight performance boosts after detoxification. Due to compute constraints, we only tested on BLOOM-7B1 model on four languages on three different tasks (see Table 7, Table 8, and Table 9) following Lai et al. (2023).
Languages Before DPO After en-DPO
vi 33.68 33.93
ru 27.46 28.14
hi 29.37 29.79
zh 37.18 37.78
Table 7: Multilingual ARC evaluation (Clark et al., 2018).
Languages Before DPO After en-DPO
vi 47.37 47.30
ru 32.60 32.84
hi 36.35 36.46
zh 50.17 50.77
Table 8: Multilingual Hellaswag evaluation (Zellers et al., 2019).
Languages Before DPO After en-DPO
vi 28.03 28.48
ru 27.09 27.59
hi 27.55 27.50
zh 29.04 29.47
Table 9: Multilingual MMLU evaluation (Hendrycks et al., 2020).
Appendix D Tradeoffs between Learning Rate, Toxicity, and Perplexity Scores
@@
Appendix E QLoRA and Multilingual Toxicity Reduction
Figure 12: Comparison between full model training and QLoRA finetuning of BLOOM-1.7B with English DPO preference tuning.
@@ Appendix E QLoRA and Multilingual Toxicity Reduction
Appendix F Bilingual Sentence Retrieval Experiment for Other LLMs
@@
Figure 13: Percentage change in expected maximum toxicity against bilingual text retrieval accuracy for BLOOM-1.7B. Correlation with Pearson-r value of 0.59 (p < 0.01)
Figure 14: Percentage change in expected maximum toxicity against bilingual text retrieval accuracy for BLOOM-7.1B. Correlation with Pearson-r value of 0.66 (p < 0.01)
Figure 15: Percentage change in expected maximum toxicity against bilingual text retrieval accuracy for Llama2-7B. Correlation with Pearson-r value of 0.78 (p < 0.01)
Appendix G Complete Table of Toxic Value Vectors
@@
Table 10: Projections of all 36 value vectors from the actual sources of toxicity - Part 1
@@
$w^{10}_{\mathrm{down},4641}$ Exclamations !", !, !".’, !», !", !, !』, !!, ?", !, !, ?".’, orrow, ?", "»., [Uncaptioned image], Who, "», +,, uu, "», Why, Your, survive, why, EEE, о , ’", Tomorrow, …
@@
$w^{21}_{\mathrm{down},7155}$ Geopolitical Tensions ίκης, リング, осс, ик, прек, ьют, [Uncaptioned image], ити, юс, енгер, бит, ета, еран, ерусалим,␣̆, олот, болгар, АТУУ, ИК, СМИ, ън, уча, американских, ек, ък, [Uncaptioned image], французского, еке
@@
$w^{11}_{\mathrm{down},3027}$ N/A OK, [Uncaptioned image], cinese,␣̆, hide, ену, хе, jade, 撲, [Uncaptioned image], sea, boys, afterwards, chines, ковой, broke, hung, енский, лё, rocks, endem, normal, ть, quit, 二世, europé, otherwise, Москва, allemande, bourg
@@
Table 11: Projections of all 36 value vectors from the actual sources of toxicity - Part 2

### v2

@@ Acknowledgements
Tài liệu
  • AI@Meta (2024) AI@Meta. 2024. Llama 3 model card.
  • Arditi et al. (2024) Andy Arditi, Oscar Obeso, Aaquib Syed, Daniel Paleka, Nina Rimsky, Wes Gurnee, and Neel Nanda. 2024. Refusal in language models is mediated by a single direction. arXiv preprint arXiv:2406.11717.
  • Artetxe and Schwenk (2019) Mikel Artetxe and Holger Schwenk. 2019. Massively multilingual sentence embeddings for zero-shot cross-lingual transfer and beyond. Transactions of the association for computational linguistics, 7:597–610.
  • Aryabumi et al. (2024) Viraat Aryabumi, John Dang, Dwarak Talupuru, Saurabh Dash, David Cairuz, Hangyu Lin, Bharat Venkitesh, Madeline Smith, Kelly Marchisio, Sebastian Ruder, Acyr Locatelli, Julia Kreutzer, Nick Frosst, Phil Blunsom, Marzieh Fadaee, Ahmet Üstün, and Sara Hooker. 2024. Aya 23: Open weight releases to further multilingual progress. Preprint, arXiv:2405.15032.
  • Bai et al. (2022) Yuntao Bai, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, Nova DasSarma, Dawn Drain, Stanislav Fort, Deep Ganguli, Tom Henighan, et al. 2022. Training a helpful and harmless assistant with reinforcement learning from human feedback. arXiv preprint arXiv:2204.05862.
  • Balestriero et al. (2023) Randall Balestriero, Romain Cosentino, and Sarath Shekkizhar. 2023. Characterizing large language model geometry solves toxicity detection and generation. arXiv preprint arXiv:2312.01648.
  • Belrose et al. (2024) Nora Belrose, David Schneider-Joseph, Shauli Ravfogel, Ryan Cotterell, Edward Raff, and Stella Biderman. 2024. Leace: Perfect linear concept erasure in closed form. Advances in Neural Information Processing Systems, 36.
  • Bereska and Gavves (2024) Leonard Bereska and Efstratios Gavves. 2024. Mechanistic interpretability for ai safety–a review. arXiv preprint arXiv:2404.14082.
  • BigScience Workshop et al. (2022) BigScience Workshop, Teven Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilić, Daniel Hesslow, Roman Castagné, Alexandra Sasha Luccioni, François Yvon, et al. 2022. Bloom: A 176b-parameter open-access multilingual language model. arXiv preprint arXiv:2211.05100.
  • cjadams et al. (2017) cjadams, Jeffrey Sorensen, Julia Elliott, Lucas Dixon, Mark McDonald, nithum, and Will Cukierski. 2017. Toxic comment classification challenge.
  • Clark et al. (2018) Peter Clark, Isaac Cowhey, Oren Etzioni, Tushar Khot, Ashish Sabharwal, Carissa Schoenick, and Oyvind Tafjord. 2018. Think you have solved question answering? try arc, the ai2 reasoning challenge. arXiv preprint arXiv:1803.05457.
  • Cui et al. (2023) Ganqu Cui, Lifan Yuan, Ning Ding, Guanming Yao, Wei Zhu, Yuan Ni, Guotong Xie, Zhiyuan Liu, and Maosong Sun. 2023. Ultrafeedback: Boosting language models with high-quality feedback. Preprint, arXiv:2310.01377.
  • de Wynter et al. (2024) Adrian de Wynter, Ishaan Watts, Nektar Ege Altıntoprak, Tua Wongsangaroonsri, Minghui Zhang, Noura Farra, Lena Baur, Samantha Claudet, Pavel Gajdusek, Can Gören, et al. 2024. Rtp-lx: Can llms evaluate toxicity in multilingual scenarios? arXiv preprint arXiv:2404.14397.
  • Dementieva et al. (2024) Daryna Dementieva, Nikolay Babakov, and Alexander Panchenko. 2024. Multiparadetox: Extending text detoxification with parallel data to new languages. arXiv preprint arXiv:2404.02037.
  • Dementieva et al. (2023) Daryna Dementieva, Daniil Moskovskiy, David Dale, and Alexander Panchenko. 2023. Exploring methods for cross-lingual text style transfer: The case of text detoxification. In Proceedings of the 13th International Joint Conference on Natural Language Processing and the 3rd Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics (Volume 1: Long Papers), pages 1083–1101, Nusa Dua, Bali. Association for Computational Linguistics.
  • Deng et al. (2024) Yue Deng, Wenxuan Zhang, Sinno Jialin Pan, and Lidong Bing. 2024. Multilingual jailbreak challenges in large language models. In The Twelfth International Conference on Learning Representations.
  • Dettmers et al. (2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. 2023. Qlora: Efficient finetuning of quantized llms. In Advances in Neural Information Processing Systems, volume 36, pages 10088–10115. Curran Associates, Inc.
  • Dufter and Schütze (2020) Philipp Dufter and Hinrich Schütze. 2020. Identifying elements essential for BERT’s multilinguality. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 4423–4437, Online. Association for Computational Linguistics.
  • Elhage et al. (2021) Nelson Elhage, Neel Nanda, Catherine Olsson, Tom Henighan, Nicholas Joseph, Ben Mann, Amanda Askell, Yuntao Bai, Anna Chen, Tom Conerly, Nova DasSarma, Dawn Drain, Deep Ganguli, Zac Hatfield-Dodds, Danny Hernandez, Andy Jones, Jackson Kernion, Liane Lovitt, Kamal Ndousse, Dario Amodei, Tom Brown, Jack Clark, Jared Kaplan, Sam McCandlish, and Chris Olah. 2021. A mathematical framework for transformer circuits. Transformer Circuits Thread. Https://transformer-circuits.pub/2021/framework/index.html.
  • Ethayarajh et al. (2024) Kawin Ethayarajh, Winnie Xu, Niklas Muennighoff, Dan Jurafsky, and Douwe Kiela. 2024. Kto: Model alignment as prospect theoretic optimization. arXiv preprint arXiv:2402.01306.
  • Fan et al. (2021) Angela Fan, Shruti Bhosale, Holger Schwenk, Zhiyi Ma, Ahmed El-Kishky, Siddharth Goyal, Mandeep Baines, Onur Celebi, Guillaume Wenzek, Vishrav Chaudhary, Naman Goyal, Tom Birch, Vitaliy Liptchinsky, Sergey Edunov, Michael Auli, and Armand Joulin. 2021. Beyond english-centric multilingual machine translation. Journal of Machine Learning Research, 22(107):1–48.
  • Ferrando et al. (2024) Javier Ferrando, Gabriele Sarti, Arianna Bisazza, and Marta R. Costa-jussà. 2024. A primer on the inner workings of transformer-based language models. Preprint, arXiv:2405.00208.
  • Gehman et al. (2020) Samuel Gehman, Suchin Gururangan, Maarten Sap, Yejin Choi, and Noah A. Smith. 2020. RealToxicityPrompts: Evaluating neural toxic degeneration in language models. In Findings of the Association for Computational Linguistics: EMNLP 2020, pages 3356–3369, Online. Association for Computational Linguistics.
  • Geva et al. (2022) Mor Geva, Avi Caciularu, Kevin Wang, and Yoav Goldberg. 2022. Transformer feed-forward layers build predictions by promoting concepts in the vocabulary space. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 30–45, Abu Dhabi, United Arab Emirates. Association for Computational Linguistics.
  • Geva et al. (2021) Mor Geva, Roei Schuster, Jonathan Berant, and Omer Levy. 2021. Transformer feed-forward layers are key-value memories. Preprint, arXiv:2012.14913.
  • Hanna et al. (2024) Michael Hanna, Ollie Liu, and Alexandre Variengien. 2024. How does gpt-2 compute greater-than?: Interpreting mathematical abilities in a pre-trained language model. Advances in Neural Information Processing Systems, 36.
  • He et al. (2016) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016. Deep residual learning for image recognition. In Proceedings of the IEEE conference on computer vision and pattern recognition, pages 770–778.
  • Hendrycks et al. (2020) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2020. Measuring massive multitask language understanding. arXiv preprint arXiv:2009.03300.
  • Holtzman et al. (2020) Ari Holtzman, Jan Buys, Li Du, Maxwell Forbes, and Yejin Choi. 2020. The curious case of neural text degeneration. In International Conference on Learning Representations.
  • Hong et al. (2024) Jiwoo Hong, Noah Lee, and James Thorne. 2024. Reference-free monolithic preference optimization with odds ratio. arXiv preprint arXiv:2403.07691.
  • Hua et al. (2024) Tianze Hua, Tian Yun, and Ellie Pavlick. 2024. mothello: When do cross-lingual representation alignment and cross-lingual transfer emerge in multilingual models? arXiv preprint arXiv:2404.12444.
  • Ivison et al. (2023) Hamish Ivison, Yizhong Wang, Valentina Pyatkin, Nathan Lambert, Matthew Peters, Pradeep Dasigi, Joel Jang, David Wadden, Noah A Smith, Iz Beltagy, et al. 2023. Camels in a changing climate: Enhancing lm adaptation with tulu 2. arXiv preprint arXiv:2311.10702.
  • Jain et al. (2024) Devansh Jain, Priyanshu Kumar, Samuel Gehman, Xuhui Zhou, Thomas Hartvigsen, and Maarten Sap. 2024. Polyglotoxicityprompts: Multilingual evaluation of neural toxic degeneration in large language models. arXiv preprint arXiv:2405.09373.
  • Khalifa et al. (2021) Muhammad Khalifa, Hady Elsahar, and Marc Dymetman. 2021. A distributional approach to controlled text generation. In International Conference on Learning Representations.
  • Kirk et al. (2024) Robert Kirk, Ishita Mediratta, Christoforos Nalmpantis, Jelena Luketina, Eric Hambro, Edward Grefenstette, and Roberta Raileanu. 2024. Understanding the effects of RLHF on LLM generalisation and diversity. In The Twelfth International Conference on Learning Representations.
  • Lai et al. (2023) Viet Lai, Chien Nguyen, Nghia Ngo, Thuat Nguyen, Franck Dernoncourt, Ryan Rossi, and Thien Nguyen. 2023. Okapi: Instruction-tuned large language models in multiple languages with reinforcement learning from human feedback. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 318–327, Singapore. Association for Computational Linguistics.
  • Lee et al. (2024) Andrew Lee, Xiaoyan Bai, Itamar Pres, Martin Wattenberg, Jonathan K Kummerfeld, and Rada Mihalcea. 2024. A mechanistic understanding of alignment algorithms: A case study on dpo and toxicity. arXiv preprint arXiv:2401.01967.
  • Lees et al. (2022) Alyssa Lees, Vinh Q Tran, Yi Tay, Jeffrey Sorensen, Jai Gupta, Donald Metzler, and Lucy Vasserman. 2022. A new generation of perspective api: Efficient multilingual character-level transformers. In Proceedings of the 28th ACM SIGKDD Conference on Knowledge Discovery and Data Mining, pages 3197–3207.
  • Liu et al. (2021) Alisa Liu, Maarten Sap, Ximing Lu, Swabha Swayamdipta, Chandra Bhagavatula, Noah A. Smith, and Yejin Choi. 2021. DExperts: Decoding-time controlled text generation with experts and anti-experts. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 6691–6706, Online. Association for Computational Linguistics.
  • Longpre et al. (2024) Shayne Longpre, Sayash Kapoor, Kevin Klyman, Ashwin Ramaswami, Rishi Bommasani, Borhane Blili-Hamelin, Yangsibo Huang, Aviya Skowron, Zheng-Xin Yong, Suhas Kotha, et al. 2024. A safe harbor for ai evaluation and red teaming. arXiv preprint arXiv:2403.04893.
  • Nanda and Bloom (2022) Neel Nanda and Joseph Bloom. 2022. Transformerlens. https://github.com/TransformerLensOrg/TransformerLens.
  • Nigatu and Raji (2024) Hellina Hailu Nigatu and Inioluwa Deborah Raji. 2024. ”i searched for a religious song in amharic and got sexual content instead”: Investigating online harm in low-resourced languages on youtube. arXiv preprint arXiv:2405.16656.
  • nostalgebraist (2020) nostalgebraist. 2020. Interpreting GPT: the logit lens. AI Alignment Forum.
  • Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022. Training language models to follow instructions with human feedback. Advances in neural information processing systems, 35:27730–27744.
  • Pozzobon et al. (2023) Luiza Pozzobon, Beyza Ermis, Patrick Lewis, and Sara Hooker. 2023. Goodtriever: Adaptive toxicity mitigation with retrieval-augmented models. In Findings of the Association for Computational Linguistics: EMNLP 2023, pages 5108–5125, Singapore. Association for Computational Linguistics.
  • Pozzobon et al. (2024) Luiza Pozzobon, Patrick Lewis, Sara Hooker, and Beyza Ermis. 2024. From one to many: Expanding the scope of toxicity mitigation in language models. arXiv preprint arXiv:2403.03893.
  • Rafailov et al. (2023) Rafael Rafailov, Archit Sharma, Eric Mitchell, Christopher D Manning, Stefano Ermon, and Chelsea Finn. 2023. Direct preference optimization: Your language model is secretly a reward model. In Advances in Neural Information Processing Systems, volume 36, pages 53728–53741. Curran Associates, Inc.
  • Raji and Dobbe (2023) Inioluwa Deborah Raji and Roel Dobbe. 2023. Concrete problems in ai safety, revisited. arXiv preprint arXiv:2401.10899.
  • Ryan et al. (2024) Michael J Ryan, William Held, and Diyi Yang. 2024. Unintended impacts of llm alignment on global representation. arXiv preprint arXiv:2402.15018.
  • Shen et al. (2024) Lingfeng Shen, Weiting Tan, Sihao Chen, Yunmo Chen, Jingyu Zhang, Haoran Xu, Boyuan Zheng, Philipp Koehn, and Daniel Khashabi. 2024. The language barrier: Dissecting safety challenges of llms in multilingual contexts. arXiv preprint arXiv:2401.13136.
  • Shliazhko et al. (2024) Oleh Shliazhko, Alena Fenogenova, Maria Tikhonova, Anastasia Kozlova, Vladislav Mikhailov, and Tatiana Shavrina. 2024. mgpt: Few-shot learners go multilingual. Transactions of the Association for Computational Linguistics, 12:58–79.
  • Touvron et al. (2023) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288.
  • Uppaal et al. (2024) Rheeya Uppaal, Apratim De, Yiting He, Yiquao Zhong, and Junjie Hu. 2024. Detox: Toxic subspace projection for model editing. arXiv preprint arXiv:2405.13967.
  • Üstün et al. (2024) Ahmet Üstün, Viraat Aryabumi, Zheng-Xin Yong, Wei-Yin Ko, Daniel D’souza, Gbemileke Onilude, Neel Bhandari, Shivalika Singh, Hui-Lee Ooi, Amr Kayid, et al. 2024. Aya model: An instruction finetuned open-access multilingual language model. arXiv preprint arXiv:2402.07827.
  • Wang et al. (2024a) Jiaan Wang, Yunlong Liang, Zengkui Sun, Yuxuan Cao, Jiarong Xu, and Fandong Meng. 2024a. Cross-lingual knowledge editing in large language models. Preprint, arXiv:2309.08952.
  • Wang et al. (2024b) Mengru Wang, Ningyu Zhang, Ziwen Xu, Zekun Xi, Shumin Deng, Yunzhi Yao, Qishen Zhang, Linyi Yang, Jindong Wang, and Huajun Chen. 2024b. Detoxifying large language models via knowledge editing. Preprint, arXiv:2403.14472.
  • Wang et al. (2023) Wenxuan Wang, Zhaopeng Tu, Chang Chen, Youliang Yuan, Jen-tse Huang, Wenxiang Jiao, and Michael R Lyu. 2023. All languages matter: On the multilingual safety of large language models. arXiv preprint arXiv:2310.00905.
  • Wei et al. (2024) Boyi Wei, Kaixuan Huang, Yangsibo Huang, Tinghao Xie, Xiangyu Qi, Mengzhou Xia, Prateek Mittal, Mengdi Wang, and Peter Henderson. 2024. Assessing the brittleness of safety alignment via pruning and low-rank modifications. arXiv preprint arXiv:2402.05162.
  • Weidinger et al. (2024) Laura Weidinger, John Mellor, Bernat Guillen Pegueroles, Nahema Marchal, Ravin Kumar, Kristian Lum, Canfer Akbulut, Mark Diaz, Stevie Bergman, Mikel Rodriguez, et al. 2024. Star: Sociotechnical approach to red teaming language models. arXiv preprint arXiv:2406.11757.
  • Wu et al. (2024) Zhaofeng Wu, Ananth Balashankar, Yoon Kim, Jacob Eisenstein, and Ahmad Beirami. 2024. Reuse your rewards: Reward model transfer for zero-shot cross-lingual alignment. arXiv preprint arXiv:2404.12318.
  • Xu et al. (2024) Haoran Xu, Amr Sharaf, Yunmo Chen, Weiting Tan, Lingfeng Shen, Benjamin Van Durme, Kenton Murray, and Young Jin Kim. 2024. Contrastive preference optimization: Pushing the boundaries of llm performance in machine translation. arXiv preprint arXiv:2401.08417.
  • Xue et al. (2021) Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2021. mT5: A massively multilingual pre-trained text-to-text transformer. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 483–498, Online. Association for Computational Linguistics.
  • Yong et al. (2023a) Zheng Xin Yong, Cristina Menghini, and Stephen Bach. 2023a. Low-resource languages jailbreak GPT-4. In Socially Responsible Language Modelling Research.
  • Yong et al. (2023b) Zheng Xin Yong, Hailey Schoelkopf, Niklas Muennighoff, Alham Fikri Aji, David Ifeoluwa Adelani, Khalid Almubarak, M Saiful Bari, Lintang Sutawika, Jungo Kasai, Ahmed Baruwa, Genta Winata, Stella Biderman, Edward Raff, Dragomir Radev, and Vassilina Nikoulina. 2023b. BLOOM+1: Adding language support to BLOOM for zero-shot prompting. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 11682–11703, Toronto, Canada. Association for Computational Linguistics.
  • Zellers et al. (2019) Rowan Zellers, Ari Holtzman, Yonatan Bisk, Ali Farhadi, and Yejin Choi. 2019. HellaSwag: Can a machine really finish your sentence? In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 4791–4800, Florence, Italy. Association for Computational Linguistics.
  • Zou et al. (2024) Andy Zou, Long Phan, Justin Wang, Derek Duenas, Maxwell Lin, Maksym Andriushchenko, Rowan Wang, Zico Kolter, Matt Fredrikson, and Dan Hendrycks. 2024. Improving alignment and robustness with short circuiting. arXiv preprint arXiv:2406.04313.
(a) Probability of generating toxic continuations
(b) Expected maximum toxicity
Hình 5: Toxicity reduction of BLOOM-1.7B BigScience Workshop et al. (2022) after DPO training.
(a) Probability of generating toxic continuations
(b) Expected maximum toxicity
Hình 6: Toxicity reduction of BLOOM-7.1B (BigScience Workshop et al., 2022) after DPO training.
(a) Probability of generating toxic continuations
(b) Expected maximum toxicity
Hình 7: Toxicity reduction of Llama2 (Touvron et al., 2023) after DPO training.
(a) Probability of generating toxic continuations
(b) Expected maximum toxicity
Hình 8: Toxicity reduction of Llama3 (AI@Meta, 2024) after DPO training.
(a) Probability of generating toxic continuations
(b) Expected maximum toxicity
Hình 9: Toxicity reduction of Aya-23 (Aryabumi et al., 2024) after DPO training.
Phụ lục A Training Details
@@ A.1 DPO Preference Tuning
Bảng 6: Hyperparameters for DPO preference tuning for mGPT and BLOOM (1.7B).
@@ A.2 Probe Training
Bảng 7: Training hyperparameters for the binary toxicity classification probe $w_{\mathrm{toxic}}$.
@@
Hình 10: Per-language perplexity distribution of mGPT continuations before and after DPO training.
Phụ lục B Distribution of Perplexity Scores
@@
Hình 11: Tradeoffs between DPO learning rate, toxicity in post-DPO generation and perplexity across 17 languages.
Phụ lục C Tradeoffs between Learning Rate, Toxicity, and Perplexity Scores
@@
Phụ lục D QLoRA and Multilingual Toxicity Reduction
Hình 12: Comparison between full model training and QLoRA finetuning of BLOOM-1.7B with English DPO preference tuning.
@@
Phụ lục E Bilingual Sentence Retrieval Experiment for Other LLMs
@@
Hình 13: Percentage change in expected maximum toxicity against bilingual text retrieval accuracy for BLOOM-1.7B. Correlation with Pearson-r value of 0.59 (p ¡ 0.01)
Hình 14: Percentage change in expected maximum toxicity against bilingual text retrieval accuracy for BLOOM-7.1B. Correlation with Pearson-r value of 0.66 (p ¡ 0.01)
Hình 15: Percentage change in expected maximum toxicity against bilingual text retrieval accuracy for Llama2-7B. Correlation with Pearson-r value of 0.78 (p ¡ 0.01)
Phụ lục F Complete Table of Toxic Value Vectors
@@
Bảng 8: Projections of all 36 value vectors from the actual sources of toxicity - Part 1
@@
$w^{10}_{\mathrm{down},4641}$ Exclamations !”, !, !”.’, !», !”, !, !』, !!, ?”, !, !, ?”.’, orrow, ?”, ”»., [Uncaptioned image], Who, ”», +,, uu, ”», Why, Your, survive, why, EEE, о , ’”, Tomorrow, …
@@
$w^{21}_{\mathrm{down},7155}$ Geopolitical Tensions ίκης, リング, осс, ик, прек, ьют, [Uncaptioned image], ити, юс, енгер, бит, ета, еран, ерусалим, ̵̆, олот, болгар, АТУУ, ИК, СМИ, ън, уча, американских, ек, ък, [Uncaptioned image], французского, еке
@@
$w^{11}_{\mathrm{down},3027}$ N/A OK, [Uncaptioned image], cinese, ̵̆, hide, ену, хе, jade, 撲, [Uncaptioned image], sea, boys, afterwards, chines, ковой, broke, hung, енский, лё, rocks, endem, normal, ть, quit, 二世, europé, otherwise, Москва, allemande, bourg
@@
Bảng 9: Projections of all 36 value vectors from the actual sources of toxicity - Part 2